40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2753     Entry time: Thu Apr 1 17:35:24 2010
Author: Koji 
Type: Update 
Category: SUS 
Subject: Working on ITMX/Y 

Steve and Koji

- We removed old ITMX/Y from the chambers. Now they are temporarily placed on the flow table at the end. Steve is looking for nice storages for the 5inch optics.

- We wiped new ITMX/Y by isopropanol as they were dusty.

- We put them into the corresponding towers. Checked the balancing and magnet arrangements with the OSEMs. They were totally fine.

- We clamped the mirrors by the EQ stops. Wrapped the towers by Al foils.

Tomorrow we will put them into the chambers.

 

Attachment 1: IMG_2353.jpg  73 kB  Uploaded Thu Apr 1 18:44:26 2010  | Hide | Hide all
IMG_2353.jpg
ELOG V3.1.3-